# Benchmarking Deep Learning and Statistical Target Detection Methods for PFM-1 Landmine Detection in UAV Hyperspectral Imagery

### v1

@@ Abstract
In recent years, unmanned aerial vehicles (UAVs) equipped with imaging sensors and automated processing algorithms have emerged as a promising tool to accelerate large-area surveys while reducing risk to human operators. Although hyperspectral imaging (HSI) enables material discrimination using spectral signatures, standardized benchmarks for UAV-based landmine detection remain scarce. In this work, we present a systematic benchmark of four classical statistical detection algorithms, including Spectral Angle Mapper (SAM), Matched Filter (MF), Adaptive Cosine Estimator (ACE), and Constrained Energy Minimization (CEM), alongside a proposed lightweight Spectral Neural Network utilizing Parametric Mish activations for PFM-1 landmine detection. We also release pixel-level binary ground truth masks (target/background) to enable standardized, reproducible evaluation. Evaluations were conducted on inert PFM-1 targets across multiple scene crops using a recently released VNIR hyperspectral dataset. Metrics such as receiver operating characteristic (ROC) curve, area under the curve (AUC), precision-recall (PR) curve, and average precision (AP) were used. While all methods achieve high ROC-AUC on an independent test set, the ACE method observes the highest AUC of 0.989. However, because target pixels are extremely sparse relative to background, ROC-AUC alone can be misleading; under precision-focused evaluation (PR and AP), the Spectral-NN outperforms classical detectors, achieving the highest AP. These results emphasize the need for precision-focused evaluation, scene-aware benchmarking, and learning-based spectral models for reliable UAV-based hyperspectral landmine detection. The code and pixel-level annotations will be released.
@@ I Introduction
This need for benchmarking is further complicated by the inherent nature of the landmine detection task. While modern transformer and attention-based architectures [4, 5] have significantly advanced hyperspectral image classification, they typically rely on abundant labeled data that are rarely met in environments where targets are extremely sparse. In contrast, classical target detection methods such as SAM [10], MF [15], ACE [9], and CEM [6] remain widely used due to interpretability and minimal training needs. However, despite their widespread use, these traditional detectors have not been systematically benchmarked on modern UAV-based hyperspectral datasets, particularly when facing the complex spectral and spatial challenges posed by PFM-1 targets.
@@ II-A Preprocessing and Ground Truth Annotation
Figure 2: Cropped regions used in this study from the original hyperspectral dataset [11]: (top) Full Region, containing all mine locations; (middle) PFM-1 Region, showing only PFM-1 targets; (bottom) Ground truth binary mask of the PFM-1 Region, with training set containing the first five PFM-1 targets (left) and the remaining two targets used for testing (right).
Seven locations containing partially or fully visible PFM-1 mines were manually annotated in ENVI to generate a binary ground truth mask (1 for target, 0 for background) resulting in 248 target pixels, using georeferenced hyperspectral and high-resolution RGB imagery for improved accuracy.
@@ II-B Statistical Target Detection Algorithms
Classical statistical target detection algorithms exploit spectral similarity and background statistics to detect a known target signature in hyperspectral imagery. Four widely used detectors, SAM [10], MF [15], ACE [9], and CEM [6], are evaluated using the ground spectral signature of PFM-1 provided by [11]. In our implementation of SAM, a negative sign was applied to the scores so that higher values correspond to stronger target responses. For consistency across MF, ACE, and CEM, the data were mean-centered, and the resulting detection scores were normalized to the range $[0,1]$ for visualization and comparison.
@@ IV Conclusion and Future Directions
Future work will investigate uncertainty-aware and spatial–spectral models to improve robustness and operational applicability. In addition, the availability of a co-registered drone-based metal detector dataset over the same test site in [12] enables exploration of multi-sensor data fusion for detecting both surface and shallowly buried PFM-1 landmines.

### v2

@@ Abstract
In recent years, unmanned aerial vehicles (UAVs) equipped with imaging sensors and automated processing algorithms have emerged as a promising tool to accelerate large-area surveys while reducing risk to human operators. Although hyperspectral imaging (HSI) enables material discrimination using spectral signatures, standardized benchmarks for UAV-based landmine detection remain scarce. In this work, we present a systematic benchmark of four classical statistical detection algorithms, including Spectral Angle Mapper (SAM), Matched Filter (MF), Adaptive Cosine Estimator (ACE), and Constrained Energy Minimization (CEM), alongside a proposed lightweight Spectral Neural Network utilizing Parametric Mish activations for PFM-1 landmine detection. We also release pixel-level binary ground truth masks (target/background) to enable standardized, reproducible evaluation. Evaluations were conducted on inert PFM-1 targets across multiple scene crops using a recently released VNIR hyperspectral dataset. Metrics such as receiver operating characteristic (ROC) curve, area under the curve (AUC), precision-recall (PR) curve, and average precision (AP) were used. While all methods achieve high ROC-AUC on an independent test set, the ACE method observes the highest AUC of 0.989. However, because target pixels are extremely sparse relative to background, ROC-AUC alone can be misleading; under precision-focused evaluation (PR and AP), the Spectral-NN outperforms classical detectors, achieving the highest AP. These results emphasize the need for precision-focused evaluation, scene-aware benchmarking, and learning-based spectral models for reliable UAV-based hyperspectral landmine detection. The code, dataset link, and pixel-level annotations are available at: https:/github.com/PrasannaPulakurthi/pfm1-hsi-benchmark.
@@ I Introduction
This need for benchmarking is further complicated by the inherent nature of the landmine detection task. While modern transformer and attention-based architectures [4, 5] have significantly advanced hyperspectral image classification, they typically rely on abundant labeled data that are rarely met in environments where targets are extremely sparse. In contrast, classical target detection methods such as SAM [12], MF [18], ACE [11, 10], and CEM [7, 6] remain widely used due to interpretability and minimal training needs. However, despite their widespread use, these traditional detectors have not been systematically benchmarked on modern UAV-based hyperspectral datasets, particularly when facing the complex spectral and spatial challenges posed by PFM-1 targets.
@@ II-A Preprocessing and Ground Truth Annotation
Figure 2: Cropped regions used in this study from the original hyperspectral dataset [14]: (top) Full Region, containing all mine locations; (middle) PFM-1 Region, showing only PFM-1 targets; (bottom) Ground truth binary mask of the PFM-1 Region, with training set containing the first five PFM-1 targets (left) and the remaining two targets used for testing (right). GSD $\approx$ 1.29 cm.
Seven locations containing partially or fully visible PFM-1 mines were manually annotated in ENVI® [13]  [^1] ENVI® is a registered trademark of NV5 Global, Inc. to generate a binary ground truth mask (1 for target, 0 for background). This resulted in 248 target pixels, using georeferenced hyperspectral and high-resolution RGB imagery for improved accuracy.
@@ II-B Statistical Target Detection Algorithms
Classical statistical target detection algorithms exploit spectral similarity and background statistics to detect a known target signature in hyperspectral imagery. Four widely used detectors, SAM [12], MF [18], ACE [11, 10], and CEM [7, 6], are evaluated using the ground spectral signature of PFM-1 provided by [14]. In our implementation of SAM, a negative sign was applied to the scores so that higher values correspond to stronger target responses. For consistency across MF, ACE, and CEM, the data were mean-centered, and the resulting detection scores were normalized to the range $[0,1]$ for visualization and comparison.
@@ IV Conclusion and Future Directions
Future work will investigate uncertainty-aware and SOTA spatial–spectral models to improve robustness and operational applicability. In addition, the availability of a co-registered drone-based metal detector dataset over the same test site in [15] enables exploration of multi-sensor data fusion for detecting both surface and shallowly buried PFM-1 landmines.